Sol Summary – December 14th

Summary Title:  Wrapping Up!

Mission Status:

Today marks the end of our scientific mission and the completion of our three remaining scientific objectives.

Our martian day sleep study ends tonight with a 4am bedtime, 12pm wake up.  Our team prepped extensively for this and arranged numerous protocols to handle crew fatigue and performance drop off due to tiredness.  Luckily we haven’t had to use any!  Our team has remained in good spirits and been able to keep to our sleep schedule well throughout our mission, with the occasional controlled nap being used to support a flagging crewmember!  We’ve focused on mentally demanding activities and music to keep the crew focused and alert in the evenings to allow us to hit our sleep targets.

Today we successfully completed our last Matryoshka objective!  This was the farthest we’ve had to travel for an EVA, which is why we left it till last to capitalize on the experience we’ve gained in the past two weeks.  We took the Deimos Rover and an ATV to ensure we had sufficient range to meet our objectives and departed from the HAB at around 1pm, arriving at our objective at around 2pm in the Blue Hills area.  Our goal was to get samples from the ridgeline as it washed out into the plane of dirt to try to identify different materials in the mixing area between the cliffs and plains.  We successfully took samples from this area and several isolated hilltops in the area and packed up from or last EVA to return home.

We weren’t going to get off easy however, as that hasn’t been our type of mission.  Due to user error when we started home the parking break was left on for ~ 5 minutes, which resulted in a precipitous power drop off from ~ 80% to 40%.  After the problem was noticed and corrected we discussed options.  The Blue Hills area is actually a fairly close walking distance from the HAB, but a long driving distance, so the decision was weather to leave the rover and walk back to the HAB or to take the long way around on the roads and trust the Rover to make it.  After assessing the fact that the Rover utilized ~ 20% battery to get to this location, had 40% remaining, and that we had a functioning ATV to get help if we broke down on the way we decided we had high confidence in getting the Rover back to the HAB and started the long drive back.  It made for an anxious drive back and frequent map checks to ensure we were going the right way but we managed to successfully coax Deimos back to the HAB just as the low battery light turned on at 20% SOC.  Another successful, and exciting, EVA!

Today we completed our Anesthesia runs ‘for score’.  Under the supervision and evaluation of our resident doctor two crewmembers at a time formed a team and had to evaluate a test mold he has created to mimic a human leg.  Inside the mold the team must use the instructional videos and procedures provided to identify various components inside the mold (veins / arteries / nerves / etc. . .), identify the proper spot to apply anesthesia, then go through the process of delivering anesthesia to the leg.  Hopefully this research and experimentation makes good progress towards developing medical trainers that can be used with astronauts on a long duration mission that may not have anyone on board with medical expertise.

Unfortunately cloud cover prevented any viewing of the Geminid meteor shower last night (of course the only cloudy night was the one with the meteor shower!)  It is currently clear and we intend to try again tonight to see if we can catch the tail end of the meteor shower.

Tomorrow we focus on MDRS cleaning, packing, and transition to the next crew.  We will also be doing an EVA strictly in the environs of MDRS to get some last filming done as requested by our journalist.

Sol Activity Summary:

  1. Matryoshka EVA Site #7a
  2. Exercise
  3. General HAB Cleaning / Maintenance / Reports
  4. Anesthesia Experiment
  5. Beginning of cleanup / packing
  6. (Hopefully)  Telescope / Geminid Observation tonight due to clouds last night!

Look Ahead Plan:

  1. EVA Filming around the HAB
  2. Solar Observation / Observatory Shutdown
  3. MDRS Cleaning
  4. Lessons Learned Compilation
  5. End of Mission Food Inventory
  6. Mission Summary
  7. Exercise
  8. General HAB Cleaning / Maintenance / Reports

Anomalies in work:

1.    HAB Water Leak continues from under EVA door.  It appears to be getting worse.

2.  Power System Misconfiguration ongoing.

Weather: Warm (above freezing)

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:



Sol Summary


Support Requested:  None

Copyright © The Mars Society. All rights reserved. | Main Site